OpenStudy (anonymous):

How do I factor 27a^3-3a

OpenStudy (anonymous):

3a(9a^2-1)

OpenStudy (anonymous):

take out the variable, then see if the constants are able to be taken out

OpenStudy (turingtest):

3a(9a^2-1)=3a(3a-1)(3a+1) difference of squares at the end

OpenStudy (anonymous):

You pull out like terms to the front: 3a(9x^2 - 1) Then you notice the second part is of the form a^2 - b^2 = (a - b) (a + b) so: 3a (3x - 1)(3x + 1)

OpenStudy (turingtest):

You pull out like terms to the front: 3a(9a^2 - 1) Then you notice the second part is of the form a^2 - b^2 = (a - b) (a + b) so: 3a (3a - 1)(3a + 1) Lance put an x in there by accident, so here is the right answer without the confusion.

OpenStudy (anonymous):

thank you I remember now